1. The following exchange rates are given:

USD 1 = AUD 1.33
USD 1 = GBP 0.56
AUD 1 = GBP 0.41

An Australian investor is looking for an arbitrage opportunity. Assuming this investor has AUD 400 available (and cannot borrow additional funds) calculate the amount of profit that can be made. Give your answer in Australian dollars and cents to the nearest cent.

Profit =

2. Before the collapse of Barings Bank in 1995 executives were thrilled when Nick Leeson announced huge arbitrage profits in trading between Singapore markets and Japanese markets. The executives even considered switching the focus of the entire bank to arbitrage. Why were the executives incredibly ignorant in not recognising a problem?

arbitrage cannot generate any profit because it has zero risk
arbitraging between two countries will not generate a profit because of the balancing effect of foreign exchange
arbitrage is risky so profits cannot continue indefinitely
arbitrage can never be substantially profitable because it has zero risk

[2  marks]

3. select a or b on the statements below giving the appropriate definition of each statement about currency trading:

To take a speculative position in a currency means: (a.) to use a trade to secure a risk free profit (b) to use a trade to deliberately create risk.

To take an arbitrage opportunity in a currency means: (a.) to use a trade to secure a risk free profit (b)to use a trade to deliberately create risk.

4. According to the law of one price:

two bonds with the same coupon rate must have the same price
a bond must have a single price in an efficient market
the price of a bond does not change between the date of purchase and maturity
the price of a bond is independent of the coupon rate
Answer #1

1. profit = $10.78

Exchange price = 400*(1/1.33)*0.56*(1/0.41) = 410.78

Current price = 400

Profit = 410.78-400 = 10.78

2. answer is option D

arbitrage can never be substantially profitable because it has zero risk

Arbitrage is used when there is no one price for identical securities in different markets and there is no risk as arbitrager trades one security for the same security in different market and thus there is less return.

3. Part 1: (b) to use a trade to deliberately create risk.

Part 2: (a.) to use a trade to secure a risk free profit

To take a speculative position in a currency means to use a trade to deliberately create risk. Speculation is associated with accepting risk deliberately.

To take an arbitrage opportunity in a currency means to use a trade to secure a risk free profit. Arbitrage is associated with exploitation of different markets to buy and sell to generate risk free profit.

4. answer is option B

a bond must have a single price in an efficient market

Identical securities must have one price in an efficient market according to the law of one price. If this law does not hold, then there exist arbitrage opportunities.